a canned juice drink is 15% ans anouther juice is 5%
How many liters of each would it take to make it 10L that is 13% juice?

Answer by stanbon(48568)

a canned juice drink is 15% ans another juice is 5%
How many liters of each would it take to make it 10L that is 13% juice?
----------------------
Equations:
Quantity Eq: canned + other = 10L
Juice Eq::: 0.15c + 0.05t = 0.13*10
-----------------------------
Substitute for c (or for t) to solve:
0.15c + (0.05)(10-c) = 0.13*10
Multiply thru by 100 to get:
15c + 5(10-c) = 13*10
15c + 50 - 5c = 130
10c = 80
c = 8L (amt. of canned juice drink needed in the mixture)
Then "other" = 10-8 = 2L (amt. of 5% juice needed in the mix)
=============================================
Cheers,
Stan H.
Reply to stanbon@comcast.net
=============================================